### v2

@@ 7 Finite-memory Nash equilibria in Büchi and co-Büchi games
This section presents our results for Büchi and co-Büchi games. For reachability, safety and shortest-path games, we have constructed finite-memory Nash equilibria from NE outcomes admitting a well-structured segment decomposition, and the resulting NEs have a memory size at most quadratic in the number of players (cf. Theorems 5.105.13 and 6.6). In Büchi and co-Büchi games, we cannot directly use the same approach to construct finite-memory NEs. Instead, we combine the classical idea of encoding an outcome to be enforced in the memory structure and decomposition-based strategies to show that from any NE in a Büchi or co-Büchi game, we can derive a finite-memory NE from the same initial state with a preferable set of winning players.
@@ 7.1 Examples
First, let us view $\mathcal{G}$ as a Büchi game. We consider the (infinite ultimately periodic) decomposition $\mathcal{D}=(v_{0}v_{1}v_{2},v_{2}v_{2},v_{2}v_{2},\ldots)$ of $\pi$. The decomposition $\mathcal{D}$ is a natural candidate from which we can build a finite-memory NE from $v_{0}$ with outcome $\pi$ with the approaches used in Theorems 5.105.13 and 6.6: each segment connects successive visits to the target of $\mathcal{P}_{2}$, whose objective is satisfied by $\pi$. However, there are no suitable NEs in which $\mathcal{P}_{2}$ follows a strategy based on $\mathcal{D}$. Indeed, if $\mathcal{P}_{2}$ uses such a strategy, then $\mathcal{P}_{1}$ can enforce their objective with an in-segment deviation from $v_{1}$ resulting in the outcome $(v_{0}v_{1})^{\omega}\in\mathsf{B\ddot{u}chi}(\{v_{1},v_{0}\})$, as $\mathcal{P}_{2}$ would not punish the deviation.
@@
Finally, we observe that the two-phase approach described above yields finite-memory strategies for which we cannot provide arena-independent memory bounds as we encode a simple history in the memory. The following example illustrates that we cannot provide arena-independent memory bounds for constrained NEs in Büchi and co-Büchi games in general.
[…]
Figure 8: The arena of Example 7.3 for $p=3$. In the game where the objective of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are respectively $\mathsf{B\ddot{u}chi}(V\setminus\{v_{4},v_{5}\})=\mathsf{coB\ddot{u}chi}(\{v_{4},v_{5}\})$ and $\mathsf{B\ddot{u}chi}(\{v_{4}\})=\mathsf{coB\ddot{u}chi}(V\setminus\{v_{4}\})$ (where $V$ is the set of vertices), $\mathcal{P}_{2}$ needs a memory of size at least $3$ in any NE from $v_{1}$ ensuring the objective of $\mathcal{P}_{2}$. Circles and squares are resp. $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ vertices.
Example 7.3.
Let $p\in\mathbb{N}\setminus\{0\}$. We define a two-player arena with $22p+$ 2 vertices such that, to obtain an NE from a given initial vertex that ensures the Büchi (and co-Büchi) objective of $\mathcal{P}_{2}$, $\mathcal{P}_{2}$ must use a strategy with at least $p$ memory states.
@@
    any strategy profile in which $\mathcal{P}_{2}$ has a finite-memory strategy of size at most $p$ is not an NE from $v_{1}$.
@@
First, we recall that $h_{q}h$ is consistent with $\sigma_{2}$ as it is a prefix of $\mathsf{Out}(\sigma,v_{1})$. The remainder of the argument is by induction. Assume that we have shown that $h_{q}h^{z}u_{0}\ldots u_{\ell+1}$ (with $z\in\mathbb{N}$ and $\ell+1\leq r$) is consistent with $\sigma_{2}$ and that $\widehat{\mathsf{up}}(h_{q}h^{z}u_{0}\ldots u_{\ell})=\widehat{\mathsf{up}}(h_{q}u_{0}\ldots u_{\ell})$. These properties hold for the base case $z=0$ and $\ell=r-1$. We show that same property holds for $\ell+1$. In the following, we abusively let $u_{r+1}$ denote $u_{0}$ to avoid treating the case $\ell+1=r$ separately.
Regarding memory states, we have that $\widehat{\mathsf{up}}(h_{q}h^{z}u_{0}\ldots u_{\ell}u_{\ell+1})=\widehat{\mathsf{up}}(h_{q}u_{0}\ldots u_{\ell}u_{\ell+1})$ directly from the induction hypothesis and the definition of $\widehat{\mathsf{up}}$. We consider two cases for the consistency. First, we assume that $u_{\ell+1}\in V_{1}$. The consistency of the relevant history follows directly by induction as $\mathcal{P}_{2}$ does not select the last transition. Second, we assume that $u_{\ell+1}\in V_{2}$. The induction hypothesis implies that $h_{q}h^{z}u_{0}\ldots u_{\ell+1}u_{\ell+2}$ is consistent with $\sigma_{2}$ if and only if $\mathsf{nxt}(\widehat{\mathsf{up}}(h_{q}h^{z}u_{0}\ldots u_{\ell}),u_{\ell+1})=u_{\ell+2}$. By consistency of $h_{q}h$ with $\sigma_{2}$, we have $\mathsf{nxt}(\widehat{\mathsf{up}}(h_{q}u_{0}\ldots u_{\ell}),u_{\ell+1})=u_{\ell+2}$, which combined with the induction hypothesis on memory updates proves the desired consistency claim.
[…]
@@ 7.2 Nash equilibria in Büchi games
In this section, we show that from any NE in a Büchi game, we can derive a finite-memory NE from the same initial vertex with the same set of satisfied objectives. As before, we build finite-memory NEs from well-structured NE outcomes obtained with a simplification process. We devise two variants of the two-phase approach described in the previous section. In Section 7.2.1, we provide a construction for NE outcomes in which some vertex occurs infinitely often. We provide a variant for other NE outcomes in Section 7.2.2.
@@ 7.2.1 Outcomes with an infinitely occurring vertex
We study NE outcomes in which a vertex occurs infinitely often. This case is a generalisation of the finite-arena case; in a finite arena, all plays contain some infinitely occurring vertex. We first show that from any NE outcome, we can derive a well-structured NE outcome satisfying the same objectives of $\mathcal{G}$. Second, we build finite-memory NEs from well-structured NE outcomes.
@@
We show that from any NE outcome $\pi^{\prime}$ in $\mathcal{G}$, we can derive an NE outcome $\pi$ with $\mathsf{first}(\pi)=\mathsf{first}(\pi^{\prime})$, $\mathsf{SatPl}_{\mathcal{G}}(\pi)=\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$ and $\pi$ admits an infinite decomposition $(\mathsf{sg}_{0},\mathsf{sg}_{1},\ldots)$ such that for all $j\geq 1$, no targets of players whose objective is not satisfied by $\pi$ occurs in $\mathsf{sg}_{j}$, $(\mathsf{sg}_{1},\mathsf{sg}_{2},\ldots)$ is periodic with period $k\in\llbracket{}n\rrbracket{}$ and either its period is one and $\mathsf{sg}_{1}$ is a simple cycle, or its period is greater and all segments $\mathsf{sg}_{1}$, $\mathsf{sg}_{2}$, …are non-trivial simple histories. Even though we obtain an infinite decomposition, its ultimate periodicity allows us to define finite-memory strategies from it.
The general idea is as follows. Let $\ell\in\mathbb{N}$ such that no targets of losing players occur in $\pi^{\prime}{\ell}$ and fix an infinitely occurring vertex $v^{\star}$ of $\pi^{\prime}$. For all $i\in\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$, fix $t_{i}\in T_{i}$ occurring after $v^{\star}$ in $\pi^{\prime}{\ell}$. For all $i\in\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$, there are histories from $v^{\star}$ to $t_{i}$ and from $t_{i}$ to $v^{\star}$ in which only vertices of $\pi^{\prime}{\ell}$ appear. It follows that, for all $i,i^{\prime}\in\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$, the vertices $t_{i}$ and $t_{i^{\prime}}$ are connected by simple histories or simple cycles consisting only of vertices from $\pi^{\prime}{\ell}$. We construct the sought decomposition by letting $\mathsf{sg}_{0}$ be a simple history from $v_{0}$ to some $t_{i}$ using vertices occurring in $\pi^{\prime}$, and then selecting the other segments among the aforementioned simple histories or cycles to obtain a periodic sequence (with period of at most $n$) such that all $t_{i}$ occur in the combination of its histories. We formalise this construction below.
@@
First, assume that $k=1$. We let $\mathsf{sg}_{0}$ be a simple history from $v_{0}$ to $t_{1}$ in which only states of $\pi$ occur ($\mathsf{sg}_{0}$ is a trivial history if $v_{0}=t_{1}$), and let $\mathsf{sg}_{1}$ be a simple cycle from $t_{1}$ to itself using only vertices occurring in $\pi{\ell}$. Let $\pi$ be the play admitting the decomposition $(\mathsf{sg}_{0},\mathsf{sg}_{1},\mathsf{sg}_{1},\ldots)$. By construction, we have $\mathsf{SatPl}_{\mathcal{G}}(\pi)=\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$, and Theorem 4.1 implies that $\pi$ is an NE outcome. Furthermore, $\pi$ and its decomposition $(\mathsf{sg}_{0},\mathsf{sg}_{1},\mathsf{sg}_{1},\ldots)$ satisfy iii and iii.
@@
We now construct finite-memory NEs from the NE outcomes given by Lemma 7.4. Assume that Lemma 7.4 provides an NE outcome $\pi$ from $v_{0}\in V$ and a decomposition $\mathcal{D}=(\mathsf{sg}_{0},\mathsf{sg}_{1},\ldots)$ of $\pi$. We implement the two-phase mechanism described in Section 7.1. For the first phase, the players enforce the history $\mathsf{sg}_{0}$. For the second phase, we switch to a strategy that is based on the periodic decomposition $\mathcal{D}^{\prime}=(\mathsf{sg}_{1},\mathsf{sg}_{2},\ldots)$. Finite memory suffices for the second phase thanks to the periodicity of $\mathcal{D}^{\prime}$: we adapt the definitions of Section 5.3 such that when reading $\mathsf{last}(\mathsf{sg}_{k})$ in memory states of the form $(\mathcal{P}_{i},k)$, we update the memory to an appropriate memory state of the form $(\mathcal{P}_{i^{\prime}},1)$.
@@
For each $i\in\llbracket{}n\rrbracket{}$, we define a Mealy machine $\mathcal{M}_{i}=(M,m_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ as follows. We define $M$ as the set $\{v_{\ell}\in V\mid 0\leq\ell\leq r\}\cup(\{\mathcal{P}_{i}\mid i\in I\}\times\llbracket{}k\rrbracket{})\cup\{\mathcal{P}_{i}\mid i\in I\}$ and let $m_{\mathsf{init}}=v_{0}$. The memory states that are vertices correspond to the first phase in the sketch of the main text, and the others to the second phase. We note that the memory bounds claimed for finite arenas follow from the simplicity of $\mathsf{sg}_{0}$ and $k\leq n$.
@@
The update function $\mathsf{up}$ is defined as follows. For all $\ell<r$, we let $\mathsf{up}(v_{\ell},v_{\ell})=v_{\ell+1}$. We let $\mathsf{up}(v_{r},v_{r})=(\mathcal{P}_{i},1)$ where $i\in I$ is such that $\mathcal{P}_{i}$ controls $v_{r}$ if $v_{r}\in V_{I}$ and $i$ is arbitrary otherwise. For all $\ell\leq r$ and $v\neq v_{\ell}$, if $\ell\geq 1$ and there exists $i\in I$ such that $v_{\ell-1}\in V_{i}$ (i.e., $v\in V_{I}$), we let $\mathsf{up}(v_{\ell},v)=\mathcal{P}_{i}$, and otherwise $\mathsf{up}(v_{\ell},v)$ is arbitrary. Intuitively, in the first phase, the Mealy machine checks that the current vertex matches the one it should be while following $\mathsf{sg}_{0}$ and switches to a special punishment state if a deviation is detected. For all states of the form $(\mathcal{P}_{i},j)\in M$ and all $v\in V$ occurring in $\mathsf{sg}_{j}$, we let $\mathsf{up}((\mathcal{P}_{i},j),v)=(\mathcal{P}_{i^{\prime}},j^{\prime})$ where (a) $\mathcal{P}_{i^{\prime}}$ is the player controlling $v$ if $v\in V_{I}$ and otherwise, $i^{\prime}=i$, and (b) $j^{\prime}=j$ if $v\neq\mathsf{last}(\mathsf{sg}_{j})$, and otherwise, if $v=\mathsf{last}(\mathsf{sg}_{j})$, we set $j^{\prime}=j+1$ if $j<k$ and $j^{\prime}=1$ otherwise. For all states of the form $(\mathcal{P}_{i},j)\in M$ and all $v\in V$ that do not occur in $\mathsf{sg}_{j}$, we let $\mathsf{up}((\mathcal{P}_{i},j),v)=\mathcal{P}_{i}$. Finally, for all $i\in I$ and $v\in V$, we let $\mathsf{up}(\mathcal{P}_{i},v)=\mathcal{P}_{i}$.
Let $i\in I$. We now define $\mathsf{nxt}_{i}$. Let $v\in V_{i}$. We first consider memory states of the form $v_{\ell}$. Fix $\ell\leq r$. If $v=v_{\ell}$, we let $\mathsf{nxt}_{i}(v_{\ell},v)=v_{\ell+1}$ if $\ell\neq r$ and let $\mathsf{nxt}_{i}(v_{r},v_{r})$ be the second vertex of $\mathsf{sg}_{1}$ (i.e., the only vertex that follows $v_{r}$ in $\mathsf{sg}_{1}$). Assume that $v\neq v_{\ell}$. If $\ell\geq 1$ and $v_{\ell-1}\in V_{I}$, we let $i^{\prime}\in I$ such that $v\in V_{i^{\prime}}$ and set $\mathsf{nxt}_{i}(v_{\ell},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$. We let $\mathsf{nxt}_{i}(v_{\ell},v)$ be arbitrary in all other cases. We now deal with memory states of the form $(\mathcal{P}_{i^{\prime}},j)$. Fix $i^{\prime}\in I$ and $j\in\llbracket{}k\rrbracket{}$. If $v$ occurs in $\mathsf{sg}_{j}$ and $v\neq\mathsf{last}(\mathsf{sg}_{j})$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)$ be the vertex following $v$ in $\mathsf{sg}_{j}$. If $v=\mathsf{last}(\mathsf{sg}_{j})$ and $j<k$ (resp. $j=k$), we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)$ be the second vertex of $\mathsf{sg}_{j+1}$ (resp. $\mathsf{sg}_{1}$). If $v$ does not occur in $\mathsf{sg}_{j}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)=\tau_{-i^{\prime}}(v)$. Finally, we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$, and otherwise we let it be arbitrary.

### v3

@@ 7 Finite-memory Nash equilibria in Büchi and co-Büchi games
This section presents our results for Büchi and co-Büchi games. For reachability, safety and shortest-path games, we have constructed finite-memory Nash equilibria from NE outcomes admitting a well-structured segment decomposition, and the resulting NEs have a memory size at most quadratic in the number of players (cf. Theorems 5.16, 5.22 and 6.6). In Büchi and co-Büchi games, we cannot directly use the same approach to construct finite-memory NEs. Instead, we combine the classical idea of encoding an outcome to be enforced in the memory structure and decomposition-based strategies to show that from any NE in a Büchi or co-Büchi game, we can derive a finite-memory NE from the same initial state with a larger set of winning players with respect to set inclusion.
@@ 7.1 Examples
First, let us view $\mathcal{G}$ as a Büchi game. We consider the (infinite ultimately periodic) decomposition $\mathcal{D}=(v_{0}v_{1}v_{2},v_{2}v_{2},v_{2}v_{2},\ldots)$ of $\pi$. The decomposition $\mathcal{D}$ is a natural candidate from which we can build a finite-memory NE from $v_{0}$ with outcome $\pi$ with the approaches used in Theorems 5.16, 5.22 and 6.6: each segment connects successive visits to the target of $\mathcal{P}_{2}$, whose objective is satisfied by $\pi$. However, there are no suitable NEs in which $\mathcal{P}_{2}$ follows a strategy based on $\mathcal{D}$. Indeed, if $\mathcal{P}_{2}$ uses such a strategy, then $\mathcal{P}_{1}$ can enforce their objective with an in-segment deviation from $v_{1}$ resulting in the outcome $(v_{0}v_{1})^{\omega}\in\mathsf{B\ddot{u}chi}(\{v_{1},v_{0}\})$, as $\mathcal{P}_{2}$ would not punish the deviation.
@@
Finally, we observe that the two-phase approach described above yields finite-memory strategies for which we cannot provide arena-independent memory bounds as we encode a simple history in the memory. The following result implies that we cannot provide arena-independent memory bounds for constrained NEs in Büchi and co-Büchi games in general.
$v_{1}$$w_{1}$$v_{2}$$w_{2}$$v_{3}$$w_{3}$$v_{4}$$v_{5}$
Figure 8: The arena used in the proof of Proposition 7.3 for $p=3$. In the game where the objective of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are respectively $\mathsf{B\ddot{u}chi}(V\setminus\{v_{4},v_{5}\})=\mathsf{coB\ddot{u}chi}(\{v_{4},v_{5}\})$ and $\mathsf{B\ddot{u}chi}(\{v_{4}\})=\mathsf{coB\ddot{u}chi}(V\setminus\{v_{4}\})$ (where $V$ is the set of vertices), $\mathcal{P}_{2}$ needs a memory of size at least $3$ in any NE from $v_{1}$ ensuring the objective of $\mathcal{P}_{2}$. Circles and squares are respectively $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ vertices.
Proposition 7.3.
Let $p\in\mathbb{N}\setminus\{0\}$. There exist a two-player game $\mathcal{G}=(\mathcal{A},(\Omega_{1},\Omega_{2}))$ where $\mathcal{A}$ is an arena with $2p+2$ vertices and $\Omega_{1}$ and $\Omega_{2}$ are both Büchi and co-Büchi objectives, and an initial vertex $v$ in $\mathcal{G}$ such that $\mathcal{P}_{2}$ requires a memory of size at least $p$ in any NE from $v$ whose outcome satisfies $\Omega_{2}$.
Proof 7.4.
@@
    any strategy profile in which $\mathcal{P}_{2}$ has a finite-memory strategy of size at most $p$ and whose outcome from $v_{1}$ satisfies $\Omega_{2}$ is not an NE from $v_{1}$.
@@
The consistency argument is by induction. We show that for all $\ell\in\mathbb{N}$, letting $a=\left\lfloor\frac{\ell}{r+1}\right\rfloor$ and $b=\ell\bmod(r+1)$, the history $h_{q}h^{a}u_{0}\ldots u_{b}$ is consistent with $\sigma_{2}$ and $\widehat{\mathsf{up}}(h_{q}h^{a}u_{0}\ldots u_{b})=\widehat{\mathsf{up}}(h_{q}u_{0}\ldots u_{b})$. For $0\leq\ell\leq r$, the former property follows from $h_{q}h$ being consistent with $\sigma_{2}$ (it is a prefix of $\mathsf{Out}(\sigma,v_{1})$) and the latter property is trivially true (because $a=0$). We now assume that the previous property holds for $\ell\in\mathbb{N}$ and show that it holds for $\ell+1$. We let $a=\left\lfloor\frac{\ell}{r+1}\right\rfloor$ and $b=\ell\bmod(r+1)$. To avoid having to distinguish the case in which $\ell+1$ is divisible by $r+1$ from the other case, we (abusively) let $u_{r+1}$ denote $u_{0}$.
Regarding memory states, we have that $\widehat{\mathsf{up}}(h_{q}h^{a}u_{0}\ldots u_{b}u_{b+1})=\widehat{\mathsf{up}}(h_{q}u_{0}\ldots u_{b}u_{b+1})$ directly from the induction hypothesis $\widehat{\mathsf{up}}(h_{q}h^{a}u_{0}\ldots u_{b})=\widehat{\mathsf{up}}(h_{q}u_{0}\ldots u_{b})$ and the definition of $\widehat{\mathsf{up}}$. We consider two cases for the consistency. First, we assume that $u_{b}\in V_{1}$. The consistency of $h_{q}h^{a}u_{0}\ldots u_{b}u_{b+1}$ with $\sigma_{2}$ follows directly by induction as $\mathcal{P}_{2}$ does not select the last transition. Second, we assume that $u_{b}\in V_{2}$. By induction, $h_{q}h^{a}u_{0}\ldots u_{b}$ is consistent with $\sigma_{2}$ and thus it remains to prove that $\sigma_{2}(h_{q}h^{a}u_{0}\ldots u_{b})=u_{b+1}$. The induction hypothesis on memory updates and the consistency of $h_{q}h$ with $\sigma_{2}$ imply that $\sigma_{2}(h_{q}h^{a}u_{0}\ldots u_{b})=\sigma_{2}(h_{q}u_{0}\ldots u_{b})=u_{b+1}$.
We have shown that in the game $\mathcal{G}$, any NE from $v_{1}$ with an outcome that is winning for $\mathcal{P}_{2}$ requires $\mathcal{P}_{2}$ to have a memory size of at least $p=\frac{|V|-2}{2}$. This shows that arena-dependent memory is required for general constrained NEs in Büchi and co-Büchi games.
@@ 7.2 Nash equilibria in Büchi games
In this section, we show that from any NE in a Büchi game, we can derive a finite-memory NE from the same initial vertex with the same set of satisfied objectives. As before, we build finite-memory NEs from well-chosen decompositions of NE outcomes obtained with a simplification process. We devise two variants of the two-phase approach described in the previous section. In Section 7.2.1, we provide a construction for NE outcomes in which some vertex occurs infinitely often. We provide a variant for other NE outcomes in Section 7.2.2.
@@ 7.2.1 Outcomes with an infinitely occurring vertex
We study NE outcomes in which a vertex occurs infinitely often. This case is a generalisation of the finite-arena case; in a finite arena, all plays contain some infinitely occurring vertex. We first show that from any NE outcome, we can derive an NE outcome satisfying the same objectives of $\mathcal{G}$ that has an ultimately periodic segment decomposition containing only simple cycles and simple histories. Second, we build finite-memory NEs from these outcomes and their decompositions.
@@
Let $\pi^{\prime}$ be an NE outcome in $\mathcal{G}$ such that some vertex occurs infinitely often in $\pi^{\prime}$. In contrast to reachability, shortest-path and safety games, we do not derive an NE outcome that has a finite simple segment decomposition from $\pi^{\prime}$. This would be too restrictive: plays that admit finite simple segment decompositions can satisfy the objectives of two players in a Büchi game only if their targets lie on the same simple cycle. Instead, we derive a lasso NE outcome $\pi$ from $\pi^{\prime}$ with the same initial vertex such that $\mathsf{SatPl}_{\mathcal{G}}(\pi)=\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$. We construct $\pi$ such that it admits an infinite decomposition $(\mathsf{sg}_{0},\mathsf{sg}_{1},\ldots)$ that satisfies three key properties that are used to implement the two-phase mechanism described in the previous section.
The prefix of $\pi$ that the players enforce in the first phase will be $\mathsf{sg}_{0}$. To ensure that no profitable deviations exist within a segment in the second phase, we ensure that for all $j\geq 1$, no targets of players whose objective is not satisfied by $\pi$ occur in $\mathsf{sg}_{j}$. For the second phase, the players use a strategy that is based on the decomposition $(\mathsf{sg}_{1},\mathsf{sg}_{2},\ldots)$. To construct a finite-memory strategy, we require that $(\mathsf{sg}_{1},\mathsf{sg}_{2},\ldots)$ is a periodic sequence with a period $k\in\llbracket{}n\rrbracket{}$ (thus $(\mathsf{sg}_{0},\mathsf{sg}_{1},\ldots)$ is ultimately periodic). Furthermore, all segments must be either simple histories or simple cycles (a decomposition-based strategy need not be well-defined otherwise). Therefore, we require that either the period is one and $\mathsf{sg}_{1}$ is a simple cycle, or its period is greater and all segments $\mathsf{sg}_{1}$, $\mathsf{sg}_{2}$, …are non-trivial simple histories.
The general idea for the derivation of $\pi$ is as follows. Let $\ell\in\mathbb{N}$ such that no targets of losing players occur in $\pi^{\prime}{\ell}$ and fix an infinitely occurring vertex $v^{\star}$ of $\pi^{\prime}$. For all $i\in\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$, fix $t_{i}\in T_{i}$ occurring after $v^{\star}$ in $\pi^{\prime}{\ell}$. For all $i\in\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$, there are histories from $v^{\star}$ to $t_{i}$ and from $t_{i}$ to $v^{\star}$ in which only vertices of $\pi^{\prime}{\ell}$ appear. It follows that, for all $i,i^{\prime}\in\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$, the vertices $t_{i}$ and $t_{i^{\prime}}$ are connected by simple histories or simple cycles consisting only of vertices from $\pi^{\prime}{\ell}$. We construct the sought decomposition by letting $\mathsf{sg}_{0}$ be a simple history from $v_{0}$ to some $t_{i}$ using vertices occurring in $\pi^{\prime}$, and then selecting the other segments among the aforementioned simple histories or cycles to obtain a periodic sequence (with period of at most $n$) such that all $t_{i}$ occur in the combination of its histories. We formalise this construction below.
@@
First, assume that $k=1$. We let $\mathsf{sg}_{0}$ be a simple history from $v_{0}$ to $t_{1}$ in which only states of $\pi$ occur ($\mathsf{sg}_{0}$ is a trivial history if $v_{0}=t_{1}$), and let $\mathsf{sg}_{1}$ be a simple cycle from $t_{1}$ to itself using only vertices occurring in $\pi{\ell}$. Let $\pi$ be the play admitting the decomposition $(\mathsf{sg}_{0},\mathsf{sg}_{1},\mathsf{sg}_{1},\ldots)$. By construction, we have $\mathsf{SatPl}_{\mathcal{G}}(\pi)=\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$, and Theorem 4.1 implies that $\pi$ is an NE outcome. Furthermore, $\pi$ and its decomposition $(\mathsf{sg}_{0},\mathsf{sg}_{1},\mathsf{sg}_{1},\ldots)$ satisfy i, ii and iii.
@@
We now construct finite-memory NEs from the NE outcomes given by Lemma 7.5. Assume that Lemma 7.5 provides an NE outcome $\pi$ from $v_{0}\in V$ and a decomposition $\mathcal{D}=(\mathsf{sg}_{0},\mathsf{sg}_{1},\ldots)$ of $\pi$. We implement the two-phase mechanism described in Section 7.1. For the first phase, the players enforce the history $\mathsf{sg}_{0}$. For the second phase, we switch to a strategy that is based on the periodic decomposition $\mathcal{D}^{\prime}=(\mathsf{sg}_{1},\mathsf{sg}_{2},\ldots)$. Finite memory suffices for the second phase thanks to the periodicity of $\mathcal{D}^{\prime}$: we adapt the updates given in Definition 5.13 such that when reading $\mathsf{last}(\mathsf{sg}_{k})$ in memory states of the form $(\mathcal{P}_{i},k)$, we update the memory to an appropriate memory state of the form $(\mathcal{P}_{i^{\prime}},1)$.
@@
For each $i\in\llbracket{}n\rrbracket{}$, we define a Mealy machine $\mathcal{M}_{i}=(M,m_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ as follows. We define $M$ as the set $\{v_{\ell}\in V\mid 0\leq\ell\leq r\}\cup(\{\mathcal{P}_{i}\mid i\in I\}\times\llbracket{}k\rrbracket{})\cup\{\mathcal{P}_{i}\mid i\in I\}$ and let $m_{\mathsf{init}}=v_{0}$. The memory states that are vertices correspond to the first phase of the informal explanation preceding the theorem, and the others to the second phase. We note that the memory bounds claimed for finite arenas follow from the simplicity of $\mathsf{sg}_{0}$ and $k\leq n$.
The update function $\mathsf{up}$ is defined as follows.
  1. (a)
    For all $\ell<r$, we let $\mathsf{up}(v_{\ell},v_{\ell})=v_{\ell+1}$.
  2. (b)
    We let $\mathsf{up}(v_{r},v_{r})=(\mathcal{P}_{i},1)$ where $i\in I$ is such that $\mathcal{P}_{i}$ controls $v_{r}$ if $v_{r}\in V_{I}$ and ii is arbitrary otherwise.
  3. (c)
    For all $\ell\leq r$ and $v\neq v_{\ell}$,
    • if $\ell\geq 1$ and there exists $i\in I$ such that $v_{\ell-1}\in V_{i}$ (in particular, $v\in V_{I}$), we let $\mathsf{up}(v_{\ell},v)=\mathcal{P}_{i}$;
    • otherwise, $\mathsf{up}(v_{\ell},v)$ is left arbitrary.
  4. (d)
    For all states of the form $(\mathcal{P}_{i},j)\in M$ and all $v\in V$ occurring in $\mathsf{sg}_{j}$, we let $\mathsf{up}((\mathcal{P}_{i},j),v)=(\mathcal{P}_{i^{\prime}},j^{\prime})$ where
    • $\mathcal{P}_{i^{\prime}}$ is the player controlling $v$ if $v\in V_{I}$ and otherwise, $i^{\prime}=i$, and
    • $j^{\prime}=j$ if $v\neq\mathsf{last}(\mathsf{sg}_{j})$, and otherwise, if $v=\mathsf{last}(\mathsf{sg}_{j})$, we set $j^{\prime}=j+1$ if $j<k$ and $j^{\prime}=1$ otherwise.
  5. (e)
    For all states of the form $(\mathcal{P}_{i},j)\in M$ and all $v\in V$ that do not occur in $\mathsf{sg}_{j}$, we let $\mathsf{up}((\mathcal{P}_{i},j),v)=\mathcal{P}_{i}$.
  6. (f)
    Finally, for all $i\in I$ and $v\in V$, we let $\mathsf{up}(\mathcal{P}_{i},v)=\mathcal{P}_{i}$.
@@
Points a, b and c relate to the first phase of our construction: the Mealy machine checks that the current vertex matches the one it should be while following $\mathsf{sg}_{0}$ and switches to a special punishment state if a deviation is detected.
Let $i\in I$. We now define $\mathsf{nxt}_{i}$. Let $v\in V_{i}$.
  1. (a)
    We first consider memory states of the form $v_{\ell}$. Fix $\ell\leq r$.
    • If $v=v_{\ell}$ and $\ell\neq r$, we let $\mathsf{nxt}_{i}(v_{\ell},v)=v_{\ell+1}$.
    • If $v=v_{\ell}$ and $\ell=r$, we let $\mathsf{nxt}_{i}(v_{r},v_{r})$ be the second vertex of $\mathsf{sg}_{1}$ (i.e., the only vertex that follows $v_{r}$ in $\mathsf{sg}_{1}$).
    • We distinguish two cases whenever $v\neq v_{\ell}$:
      • if $\ell\geq 1$ and there exists $i^{\prime}\in I$ such that $v_{\ell-1}\in V_{i^{\prime}}$ and $i^{\prime}\neq i$, we let $\mathsf{nxt}_{i}(v_{\ell},v)=\tau_{-i^{\prime}}(v)$;
      • we let $\mathsf{nxt}_{i}(v_{\ell},v)$ be arbitrary in all other cases.
  2. (b)
    We now deal with memory states of the form $(\mathcal{P}_{i^{\prime}},j)$. Fix $i^{\prime}\in I$ and $j\in\llbracket{}k\rrbracket{}$.
    • If $v$ occurs in $\mathsf{sg}_{j}$ and $v\neq\mathsf{last}(\mathsf{sg}_{j})$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)$ be the vertex following $v$ in $\mathsf{sg}_{j}$.
    • If $v=\mathsf{last}(\mathsf{sg}_{j})$ and $j<k$ (resp. $j=k$), we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)$ be the second vertex of $\mathsf{sg}_{j+1}$ (resp. $\mathsf{sg}_{1}$).
    • If $v$ does not occur in $\mathsf{sg}_{j}$ and $i\neq i^{\prime}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)=\tau_{-i^{\prime}}(v)$.
    • If $v$ does not occur in $\mathsf{sg}_{j}$ and $i=i^{\prime}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},j),v)$ be arbitrary.
  3. (c)
    Finally, we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$, and otherwise we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)$ be arbitrary.